OpenStudy (anonymous):

For which value of x does the graph of y = 3x^2 - 2x - 5 cross the x-axis?

OpenStudy (espex):

Set the equation equal to zero, factor, solve for x.

OpenStudy (anonymous):

3x^2 − 2x − 5=0 (x+1)(3x-5)=0 x+1=0 ,3x-5=0 x=-1 ,3x=5 x=5/3 so x=-1 and x=5/3 Answer x=-1 Hope this helps!